Suppose that I will pay you $8.00 for every problem in this chapter that you solve correctly but you have to pay me $5.00 for every problem done incorrectly. If at the end of the 26 problems neither of us owe each other any money, how many problems did you solve correctly?
Answer by mananth(16949)

correct problem $8
wrong problem $-5
let correct problems be x
wrong problems will be 26-x
8x+-5(26-x)=0
8x-130+5x= 0
13x=130
/13
x= 10
10 correct & 16 wrong